# Reliable Grid Forecasting: State Space Models for Safety-Critical Energy Systems

### v3

@@ Abstract
Accurate grid load forecasting is safety-critical: under-predictions risk supply shortfalls, while symmetric error metrics mask this operational asymmetry. We introduce a grid-specific evaluation framework—Asymmetric MAPE, Under-Prediction Rate, and Reserve Margin—that directly measures operational risk rather than statistical accuracy alone.
Using this framework, we conduct a systematic evaluation of Mamba-based State Space Models for California grid forecasting on a weather-aligned CA ISO-TAC dataset spanning Nov 2023–Nov 2025 (84,498 hourly records across 5 transmission areas). Our analysis reveals that standard accuracy metrics are poor proxies for operational safety: models with identical MAPE can require vastly different reserve margins.
We demonstrate that forecast errors are weakly but statistically significantly associated with temperature (), motivating weather-aware modeling rather than loss function modification alone. The S-Mamba model achieves the lowest Reserve margin (14.12%) compared to 16.66% for iTransformer, demonstrating superior forecast reliability under a 99.5th-percentile tail-risk reserve proxy.
@@
California’s grid exemplifies the challenges facing high-renewable systems. Solar and wind now account for 25% of generation [2], with behind-the-meter (BTM) solar contributing an additional 10–15% of midday native load. This invisible generation creates the “duck curve”—deep midday net load trough followed by steep evening ramps—that evolves annually as BTM capacity expands. Compounding this non-stationarity, climate-driven extreme weather events increasingly trigger non-linear demand spikes that defy historical patterns [3].
@@
We present a systematic evaluation of Mamba architectures for reliable California grid load forecasting. We benchmark against iTransformer and the foundation model Chronos, but shift the evaluation focus from pure accuracy to operational safety. Our analysis reveals that standard training objectives fail to capture the risk profile needed by operators; S-Mamba achieves a Reserve margin of 14.12%, significantly lower than the 16.66% required by iTransformer. Furthermore, we demonstrate that explicit weather integration is critical for safety, reducing the frequency of extreme (1000 MW) error events by 40%.
@@
Grid-Specific Evaluation Framework. We formalize operational risk metrics (-MAPE, Under-Prediction Rate, Reserve Margin) that capture the asymmetric costs of forecast errors, demonstrating that standard metrics are poor proxies for grid safety (Section 2.3).
@@
Weather Integration for SSMs. We develop and systematically evaluate thermal-lag-aligned weather fusion strategies for Mamba architectures, achieving 40% reduction in extreme errors (Section 6.5).
@@
Loss vs. Safety Analysis. We show that loss reweighting alone cannot substitute for covariate modeling: quantile loss reduces 1000 MW errors by only 1.6% on CA ISO-TAC test windows, while weather-aware modeling yields materially larger tail-risk improvements (Section 6).
@@ 2.3 Grid-Specific Evaluation Metrics
2.3.1 The Asymmetry of Grid Operations
@@ 2.3.2 Proposed Metrics
- Asymmetric MAPE (-MAPE).
We define a weighted MAPE that applies a penalty factor to under-predictions:
(5) In this work, we use , reflecting the industry consensus that under-prediction risks are approximately double the cost of over-prediction for day-ahead planning.
@@
- Reserve Margin Required.
@@
The additional upward capacity required to cover 99.5% of the under-forecast error distribution. We use the 99.5th percentile as a practical tail-risk proxy (motivated by resource adequacy practice; e.g., [11]). We report this both in megawatts (MW) and as a percentage of the point forecast (scale-free):
@@
(7) (8) Adding to the point forecast covers 99.5% of historical under-prediction events at the evaluated horizon. We report both MW and percentage: MW is normalization-free, while the percentage form is directly interpretable as an add-on to the scheduled point forecast. Because percentage-normalized reserve can be biased by systematic forecast inflation, we always interpret Reserve jointly with -MAPE and UPR to ensure improvements do not come from trivial over-forecasting.
2.3.3 Metrics Matter: A Case Study
To illustrate the operational reality of these metrics, we compare two architectures from our experimental results (Table 2) applied to a hypothetical 25,000 MW California grid system.
- PowerMamba:
Achieves better average accuracy (4.37% MAPE) but a high Under-Prediction Rate (58.8%) and Reserve Requirement (21.68%).
- S-Mamba:
Achieves slightly worse accuracy (4.77% MAPE) but significantly lower UPR (40.9%) and Reserve Requirement (14.12%).
Operational Implication.
Although PowerMamba appears superior by only looking at MAPE, its error structure is operationally expensive. The Reserve difference () implies that, for a representative 25,000 MW day-ahead point forecast, an operator using PowerMamba must schedule an additional 1,890 MW of spinning reserves every hour ().
Assuming a typical reserve holding cost of $10/MW-h, deploying the “more accurate” PowerMamba model would incur approximately $453,600 per day in excess ancillary service costs compared to S-Mamba. This paradox—where the model with better MAPE costs more to operate—demonstrates why our evaluation prioritizes -MAPE, UPR, and Reserve metrics alongside standard accuracy.
2.3.4 From Metrics to Optimization
While -MAPE and UPR are excellent for evaluation, they are non-differentiable and difficult to optimize directly. To bridge this gap, we employ Quantile Loss (Pinball Loss) as a differentiable surrogate during the training process (specifically Phase 3, see Section 5.5).
The quantile loss for a target quantile is defined as:
By training on high quantiles (e.g., ), we effectively apply a higher gradient penalty to under-predictions (weighted by ) than over-predictions (weighted by ). This aligns the model’s statistical objective with the operator’s economic objective, naturally suppressing the UPR and improving -MAPE without requiring custom non-convex loss functions. We evaluate empirically whether this theoretical alignment translates to practice in Section 6.
2.4 CAISO Market Context and Benchmarks
The California Independent System Operator (CAISO) operates two primary energy markets demanding distinct forecast horizons: the Day-Ahead Market (DAM) requiring hourly forecasts submitted by 10:00 AM the prior day, and the Real-Time Market (RTM) requiring 5-minute granularity. A defining feature of the CAISO control area is the massive penetration of Behind-the-Meter (BTM) solar, which serves approximately 10–15% of native load during midday.
This BTM capacity physically disconnects “gross load” (total consumption) from “net load” (visible grid demand), creating the well-known “duck curve” phenomenon. This introduces a critical weather dependency: cloud cover variations cause instantaneous ramp events that statistical baselines struggle to predict.
Operational load forecasts serve as the industry benchmark. During the July 2024 heat wave (July 4–12), third-party analysis reported CAISO’s day-ahead forecast achieved 4.55% MAPE, while commercial forecasting services achieved 2.65% MAPE—a 41% improvement [12]. These benchmarks establish minimum performance targets for operational deployment.
Accurate forecasting during extreme weather events is critical. The July 2024 analysis revealed CAISO’s peak load forecast error of 3,211 MW during the heat wave, highlighting the importance of weather-integrated forecasting approaches.
@@ 3.1 Deep Learning for Time Series
Linear Models. Recent work has challenged the necessity of deep learning for time series forecasting. DLinear [15] demonstrates that simple linear models often match or exceed Transformer performance on standard benchmarks, raising questions about architectural complexity. However, linear models cannot capture non-linear temperature-load relationships that dominate during extreme weather events—precisely when forecast accuracy is most critical for grid operations. Our evaluation focuses on models capable of learning these non-linear interactions.
@@ 3.3 Weather-Integrated Forecasting
The relationship between weather and electricity demand is well-established [10]. Temperature-based load models form the basis of many operational forecasting systems, with heating and cooling degree days serving as primary predictors. However, integrating weather into deep learning architectures remains challenging due to temporal misalignment: building thermal mass introduces response lags of 2–6 hours between temperature changes and load response [17]. Prior work has addressed this through feature engineering [18] or attention mechanisms [9], but systematic evaluation of weather integration strategies for state space models is lacking.
@@ 4.1.2 PowerMamba: Physics-Aware Decomposition
4.1.3 Mamba-ProbTSF: From Point Signals to Risk Management
@@ 5.1 Dataset
5.2 Thermal Lag Analysis and Justification
@@
A critical design decision for weather-integrated forecasting is the temporal alignment between meteorological variables and load response. Buildings do not respond instantaneously to temperature changes; rather, the thermal mass of walls, floors, and furnishings creates a delayed response in HVAC demand. We conducted a systematic cross-correlation analysis to determine optimal lag values empirically.
@@
Cross-Correlation Methodology. For each weather covariate and the load signal , we computed the Pearson correlation coefficient at lag hours:
The optimal lag .
Empirical Results. Analysis of the Nov 2023–Nov 2025 hourly dataset (84,498 records across 5 TAC areas) reveals distinct lag patterns by covariate type:
Temperature (dry bulb): Peak correlation at hours (), consistent with commercial building thermal time constants of 2–4 hours reported in building physics literature [17]. Residential buildings show shorter lags (1–2 hours) while large commercial buildings exhibit lags up to 6 hours.
Solar radiation (GHI): Peak correlation at hour (), reflecting the more immediate effect of solar heat gain through windows and the subsequent building envelope heating.
Humidity: Peak correlation at hours (), as latent cooling loads follow similar thermal dynamics to sensible loads.
Wind speed: Minimal lag (– hour, ), as ventilation effects manifest more quickly.
Physical Interpretation. The observed 2–6 hour lag range aligns with building thermal dynamics theory. Seem [17] models commercial building response using first-order thermal networks with time constants , where is thermal resistance and is thermal capacitance. For typical California commercial buildings:
Light construction (steel frame, single-pane): – hours
Medium construction (concrete tilt-up): – hours
Heavy construction (concrete/masonry): – hours
The aggregate California load reflects a mixture of building types, with commercial buildings dominating peak demand. Our empirically-derived lags of 2–4 hours for temperature-related covariates are consistent with this physical understanding.
@@
Horizon-Adaptive Alignment. For multi-horizon forecasting, we apply horizon-adaptive lag adjustment. When forecasting load at time for horizon , the effective lag becomes . Weather features at time are aligned to predict load response at , ensuring consistent temporal relationships across all forecast horizons.
Train/Validation/Test Split. We employ chronological splitting to respect temporal dependencies: 70% training, 15% validation, and 15% test within each TAC-area time series. Due to minor missingness differences by region, exact split dates vary slightly by utility; for CA ISO-TAC, the split is train (2023-11-01–2025-04-09), validation (2025-04-09–2025-07-31), and test (2025-07-31–2025-11-22). This non-overlapping split ensures evaluation on future, unseen data—critical for operational deployment assessment. For evaluation, we construct sliding windows with 240-hour context and 48-hour prediction horizon; for CA ISO-TAC, the 2,649-hour test segment yields 2,362 forecast windows.

### v4

# Beyond MAPE: Tail-Risk Metrics and Bias Control for Reliable Grid Forecasting

@@ Abstract
Accurate grid load forecasting is safety-critical: under-predictions risk supply shortfalls, while symmetric error metrics can mask this operational asymmetry. We introduce an operator-legible evaluation framework—Under-Prediction Rate (UPR), tail Reserve requirements, and explicit inflation diagnostics (Bias24h/OPR)—to quantify one-sided reliability risk beyond MAPE.
Using this framework, we evaluate state space models (Mamba variants) and strong baselines on a weather-aligned California Independent System Operator (CAISO) dataset spanning Nov 2023–Nov 2025 (84,498 hourly records across 5 regional transmission areas) under a rolling-origin walk-forward backtest. We develop and evaluate thermal-lag-aligned weather fusion strategies for these architectures.
Our results demonstrate that standard accuracy metrics are insufficient proxies for operational safety: models with comparable MAPE can imply materially different tail reserve requirements (Reserve). We show that explicit weather integration narrows error distributions, reducing the impact of temperature-driven demand spikes. Furthermore, while probabilistic calibration reduces large-error events, it can induce systematic schedule inflation. We introduce Bias/OPR-constrained objectives to enable auditable trade-offs between minimizing tail risk and preventing trivial over-forecasting.
@@
California’s grid exemplifies the challenges facing high-renewable systems. Utility-scale solar and wind contribute roughly one quarter of CAISO system energy [2], with behind-the-meter (BTM) solar capacity exceeding 17 GW in California [3]. This invisible generation creates the “duck curve”—deep midday net load trough followed by steep evening ramps—that evolves annually as BTM capacity expands. Compounding this non-stationarity, climate-driven extreme weather events increasingly trigger non-linear demand spikes that defy historical patterns [4].
@@
We present a systematic evaluation of Mamba architectures for reliable California grid load forecasting. Benchmarking against iTransformer and the foundation model Chronos, we shift the evaluation focus from pure accuracy to operational safety. Our walk-forward results highlight that accuracy alone is not a sufficient proxy for operational risk: Reserve can vary substantially across models even when MAPE is comparable. Furthermore, we demonstrate that explicit weather integration is critical for safety by narrowing error distributions and reducing extreme error events (Section 6.3).
@@
Grid-Specific Evaluation Framework. We formalize operational risk metrics (Under-Prediction Rate, tail Reserve requirements, Bias/OPR diagnostics) that capture asymmetric operational costs and prevent “fake safety” via systematic forecast inflation (Section 2.3).
@@
Loss vs. Safety Analysis. We show that proper probabilistic calibration (multi-quantile pinball) can reduce large-error events on fixed evaluation splits, but can also introduce systematic inflation unless Bias/OPR are explicitly constrained and jointly reported with tail-risk metrics (Section 6).
@@ 2.3 Grid-Specific Evaluation Metrics
2.3.1 Operational Cost Asymmetry
@@ 2.3.2 Proposed Metrics
- Over-Prediction Rate (OPR).
The frequency of over-forecasting events exposes systematic inflation in the scheduled (median) forecast:
(6) In safety-critical settings, OPR provides a simple operator-legible diagnostic for whether improvements in one-sided tail risk are achieved by trivial over-scheduling.
- Upward Reserve Requirement (Reserve).
@@
The additional upward capacity required to cover 99.5% of the under-forecast error distribution. We use the 99.5th percentile as a practical tail-risk proxy motivated by probabilistic adequacy practice (e.g., LOLE-style planning criteria; [11]). We report this both in megawatts (MW) and as a percentage of the point forecast (scale-free):
@@
(7) (8) Adding to the point forecast covers 99.5% of historical under-prediction events at the evaluated horizon. We report both MW and percentage: MW is normalization-free, while the percentage form is directly interpretable as an add-on to the scheduled point forecast. Because percentage-normalized reserve can be biased by systematic forecast inflation, we interpret Reserve jointly with UPR and explicit bias diagnostics (Bias24h / OPR) to ensure improvements do not come from trivial over-forecasting.
2.3.3 Motivating Example
To illustrate the operational reality of these metrics, we examine empirical results (detailed later in Table 2) where tail-risk metrics and bias diagnostics move in different directions. For example, a model trained with a tail-focused loss might ostensibly reduce reserve requirements (Reserve), but do so by simply inflating the entire forecast—increasing Bias24h and Over-Prediction Rate (OPR) to unacceptable levels. This motivates our core reporting principle: any improvement in one-sided risk metrics must be interpreted jointly with bias diagnostics to avoid “fake safety” via systematic inflation.
2.3.4 Differentiable Surrogates for Risk Metrics
While UPR and tail reserve requirements are excellent for evaluation, they are non-differentiable and difficult to optimize directly. To bridge this gap, we employ a Bias-Constrained Probabilistic Objective (detailed in Section 4.3) as a differentiable surrogate during training. This objective combines a weighted multi-quantile pinball loss to learn a calibrated distribution with explicit hinge penalties on bias and over-prediction rate to prevent systematic forecast inflation.
In principle, emphasizing upper-tail quantiles increases the gradient penalty on under-predictions relative to over-predictions, aligning learning with asymmetric operational costs. However, as we demonstrate in Section 6, single-quantile training can be “gamed”; our constrained formulation (Section 4.3) ensures that improvements in tail risk do not come at the expense of acceptable schedule bias.
2.4 Value-Oriented Learning
Traditional forecasting objectives (MSE/MAPE) operate under the assumption that prediction accuracy proxies decision quality. However, in grid operations, the cost function is highly asymmetric and constrained. The “Value-Oriented” or “Smart Predict-then-Optimize” paradigm [12] challenges the separation of prediction and downstream optimization, proposing that models should be trained to minimize the final operational cost. While theoretically optimal, full end-to-end bilevel optimization is often computationally intractable for large-scale deep learning. In this work, we adopt a scalable realization of this philosophy: instead of embedding the full extensive-form optimization solver, we encode the operational value function directly into a bias-constrained probabilistic objective (Section 4.3).
@@ 3.1 Deep Learning for Time Series
Linear Models. Recent work has challenged the necessity of deep learning for time series forecasting. DLinear [15] demonstrates that simple linear models often match or exceed Transformer performance on standard benchmarks, raising questions about architectural complexity. However, linear models cannot capture non-linear temperature-load relationships that dominate during extreme weather events—precisely when forecast accuracy is most critical for grid operations. While recent “value-oriented” and end-to-end approaches propose embedding optimization layers directly into the training loop [12], these methods can be computationally prohibitive for large-scale deep learning sequence models. We instead propose a bias-constrained objective as a scalable, differentiable surrogate to align training with asymmetric operational costs. Our evaluation focuses on models capable of learning these non-linear interactions while maintaining computational tractability.
@@ 3.2 State Space Models for Time Series
3.3 End-to-End and Value-Oriented Learning
Recent advancements in "Smart Predict-then-Optimize" (SPO) have challenged the traditional separation of forecasting and decision-making. Value-oriented approaches [12] propose bi-level optimization frameworks where the forecasting model is updated based on the quality of downstream decisions (e.g., unit commitment costs) rather than proxy metrics like MSE. Similarly, end-to-end structural learning [17] embeds differentiable optimization layers to disentangle unobservable components (like baseline demand vs. price response) from aggregate signals.
Our work builds on these philosophies but focuses on scalability. Instead of solving extensive optimization problems within the training loop—which is often prohibitive for large-scale transformers—we propose a bias-constrained probabilistic objective as a differentiable surrogate that aligns the model with asymmetric operational values while maintaining the efficiency of standard backpropagation.
@@ 3.4 Weather-Integrated Forecasting
The relationship between weather and electricity demand is well-established [1]. Temperature-based load models form the basis of many operational forecasting systems, with heating and cooling degree days serving as primary predictors. However, integrating weather into deep learning architectures remains challenging due to temporal misalignment: building thermal mass introduces response lags of 2–6 hours between temperature changes and load response [18]. Prior work has addressed this through feature engineering and attention-based fusion strategies [5], but systematic evaluation of weather integration strategies for state space models is lacking.
@@ 4.1.2 PowerMamba: Physics-Aware Decomposition
4.1.3 Mamba-ProbTSF: Probabilistic Output Heads
@@
Alignment with our evaluations. In our main walk-forward results we evaluate this Gaussian-head variant as a probabilistic baseline; in the fixed-split loss-function ablations we fine-tune weather-initialized checkpoints with multi-quantile heads to study calibration and Bias/OPR-constrained objectives (Section 4.3, Table 2).
@@
4.2 Weather-Integrated Architectures
Motivated by the dependency of load on weather (analyzed in Section 6), we developed weather-integrated variants of each architecture. The key challenge is fusing meteorological covariates (temperature, humidity, solar radiation) with the load signal while respecting building thermal lag—HVAC systems respond to temperature changes with delays of 2–6 hours depending on building mass.
4.3 Bias-Constrained Probabilistic Objective
Grid operations induce an asymmetric cost for forecast errors: under-predictions can require expensive upward balancing and reserves, while over-predictions typically incur lower (but non-zero) costs from over-commit, curtailment, and redispatch. We represent this asymmetry using a cost ratio
where and are marginal costs per MWh associated with under- and over-forecasting, respectively. In practice, includes reliability and scarcity costs that are not fully captured by average market spreads. We therefore use CAISO market data to estimate a bias-independent price-spread asymmetry anchor (Appendix F), and map it to an operational stance via an explicit reliability premium :
This makes the risk-aversion choice auditable: grounds asymmetry empirically, while documents the operator policy preference.
From cost asymmetry to an operational target quantile. Under a piecewise-linear asymmetric cost model, the Bayes-optimal quantile is
We set a conservative risk-averse training quantile by combining the market anchor with the operator stance:
ensuring we never train below the median when the objective is reliability-seeking. (Appendix F reports that can imply in calendar-year averages; we interpret this as evidence that market spreads alone are not a complete proxy for reliability cost, motivating the explicit premium.)
Proper probabilistic training (multi-quantile pinball). Rather than training a single high quantile (which can induce systematic forecast inflation), we train a calibrated predictive distribution using multiple quantiles and a weighted proper scoring rule. For each , define the pinball loss
and the weighted multi-quantile objective over horizon :
This objective is strictly proper for quantile forecasts; when is dense, it approximates distributional scores such as CRPS. Multi-quantile pinball is also the standard evaluation objective in benchmark competitions for probabilistic load forecasting (e.g., GEFCom), making it familiar and auditable for practitioners.
Why single-quantile pinball can be “gamed.” Training only a high quantile (e.g., ) can reduce under-prediction frequency by shifting the conditional location upward. This may improve one-sided risk metrics (e.g., UPR, Reserve) while introducing systematic positive bias that is operationally unacceptable (inflating schedules and masking true uncertainty). We therefore couple probabilistic training with explicit bias controls and report bias diagnostics alongside risk metrics.
Bias constraint to avoid trivial “safety” via over-forecasting. To prevent degeneracy via systematic inflation, we add an explicit positive-bias constraint on the median (scheduled) forecast at a target horizon (e.g., 24h). Let be the index for the target horizon (24h ), and define the mean bias at that horizon using the median quantile :
We implement a hinge-penalized constraint:
where discourages systematic over-forecast (inflation) and is an auditable bias budget (we use by default).
OPR-style constraint (frequency of over-forecasting). Bias controls the mean shift but does not directly bound the frequency of over-predictions. We therefore optionally constrain the over-prediction rate at horizon by penalizing violations of an OPR budget :
approximating the indicator smoothly with a sigmoid temperature :
and adding a hinge penalty:
Practically, we use these constrained objectives during fine-tuning to learn a calibrated distribution while preventing trivial “safety” via inflation. We then use (via ) to select an operationally conservative schedule or reserve policy from the learned distribution, and we report Bias/OPR jointly with tail-risk metrics to ensure improvements are not attributable to trivial inflation. Appendix F describes how we empirically estimate from CAISO market data [20].
@@ 5.1 Dataset
5.2 CAISO Market Context and Benchmarks
CAISO operations place a premium on day-ahead reliability: the Day-Ahead Market (DAM) requires hourly schedules submitted by 10:00 AM the prior day, while the Real-Time Market (RTM) clears at 5-minute granularity. A defining feature of the CAISO region is the rapid growth of behind-the-meter (BTM) solar; California BTM solar capacity exceeds 17 GW [3]. This “invisible” generation deepens the midday net-load trough and steepens evening ramps (the “duck curve”), increasing the importance of weather-aware forecasting.
Operational forecast performance provides a practical benchmark. During the July 2024 heat wave (July 4–12), third-party analysis reported CAISO’s day-ahead forecast achieved 4.55% MAPE, while commercial forecasting services achieved 2.65% MAPE (41% improvement) [21]. The same analysis reported a CAISO peak load forecast error of 3,211 MW during the event, highlighting the operational importance of reducing tail errors during extreme weather.
5.3 Thermal Lag Parameter Determination
@@
A critical design decision for weather-integrated forecasting is the temporal alignment between meteorological variables and load response. Buildings do not respond instantaneously to temperature changes; rather, the thermal mass of walls, floors, and furnishings creates a delayed response in HVAC demand.
@@
Cross-Correlation Analysis. To determine the optimal lag parameters empirically, we computed the Pearson correlation coefficient between each weather covariate and load at lags hours:
We selected the optimal lag for each covariate.
Identified Lag Parameters. Analysis of the dataset reveals distinct response patterns consistent with building physics [18]:
Temperature (dry bulb): hours (), reflecting the thermal time constants of commercial building envelopes (typically 2–4 hours).
Solar radiation (GHI): hour (), corresponding to rapid solar heat gain through fenestration.
Humidity: hours (), tracking latent cooling load dynamics.
Wind speed: – hour (), indicating immediate ventilation effects.
@@
Horizon-Adaptive Alignment. These empirically determined lags are applied during feature construction. When forecasting load at time for horizon , we align weather features from time to predict load response at , ensuring consistent temporal causality across all forecast horizons.
Evaluation Protocol (Walk-Forward Backtest). We evaluate using rolling-origin walk-forward backtesting over Nov 2023–Nov 2025. Starting with an initial 180-day training history, we refit every 90 days using an expanding training window; each refit uses a 30-day validation window immediately preceding the cutoff for early stopping. We then evaluate one forecast per day (stride 24h) over the subsequent block until the next cutoff (48h horizon). All reported metrics in Tables 1–2 are computed on the concatenated set of walk-forward timestamps (477 evaluation windows for CA ISO-TAC).